# Growth problems for representations of finite monoids

### v1

@@ 4. Examples
4A. Full transformation monoids $T_{n}$, and symmetric inverse monoids $I_{n}$
Let $T_{n}$ and $I_{n}$ denote the full transformation monoid and the symmetric inverse monoid on $n$ elements, respectively. In both cases the group of units is $G=S_{n}$.
[…]
@@
(4A.1) $b(n)\sim a(n)=\sum_{z=0}^{\left\lfloor m/2\right\rfloor}\frac{1}{(m-2z)!z!2^{z}}\cdot(\dim V)^{n}=k(n)\sim l(n).$
@@
The statements follow from Section 3B when $M=I_{n}$ as $kI_{n}$ is semisimple under the assumptions of the theorem (see [Ste16b, Corollary 9.4]). By [Ste16a, Corollary A.3], if $W$ is any $kS_{n}$-module which is not the sign representation, then $\operatorname{Ind}_{G}(W)$ is injective as a $kT_{n}$-module. Thus the assumptions of Section 3A are satisfied for $T_{n}$ under our assumptions, and the asymptotic formula for $S_{n}$ given in [CEO24, Example 2.3] now implies the result for $a(n)$. The statement about $k(n)$ follows from Section 3B. ∎
[…]
One might expect that for $T_{n}$ we can also weaken the requirement that $p=0$ to $p\nmid n!$. However, this is nontrivial to show, as monoid algebras need not have the same structure as we vary between characteristics that do not divide the order of any maximal subgroup. However, for $p$ sufficiently large the result of Section 4A will hold. $\Diamond$
Since $kI_{n}$ is semisimple when $p\nmid n!$, the exact formula for $b(n)=l(n)$ can be calculated from Section 3B. For example, if $V$ is the irreducible $\mathbb{C}I_{3}$-module with character values $0,1,2,0,3,1,0$, then using the character table and decomposition matrix for $\mathbb{C}I_{3}$ which can be found in [Ste16b, Example 9.17], in this case we obtain
@@
We plot below in Figure 2 the fusion graph and action matrix for the 12-dimensional projective $kT_{4}$-module in characteristic 0, and in Figure 3 give the ratio $b(n)/a(n)$ in this case. In this case $\lambda^{\textrm{sec}}=2$, so $\lambda^{\textrm{sec}/\dim V}=1/6$, which explains the rapid convergence. As usual, we approximated the characteristic zero behavior by doing the computation in a large enough finite field.
@@
4B. Matrix monoids $\operatorname{M}(2,q)$
@@
[…]
@@
(4B.1) $a(n)=\begin{cases}\frac{q^{2}+(-1)^{n}}{(q+1)(q^{2}-q)}\cdot(\dim V)^{n}&\text{if $\epsilon$ acts as scalar on $\operatorname{Res}_{G}(V)$},\vspace{5pt}\\ \frac{q^{2}}{(q+1)(q^{2}-q)}\cdot(\dim V)^{n}&\text{else}.\end{cases}$
@@
Since $\operatorname{M}(2,q)$ is semisimple, by Section 3A (or Section 3B) in this case $a(n)$ is determined by the character table of $G$, which can be found in e.g. [FH91, §5.2]. The sum over the dimensions of irreducible $kG$-modules is $q^{2}(q-1)$, and when $q$ is odd the sum over the column of the character table corresponding to $\epsilon$ is $q-1$. The sum over all other columns in the center vanishes, yielding our formula. ∎
@@
[…]
@@
(4B.2) $a(n)=\begin{cases}\frac{(p+1)^{r}}{2^{r}(q+1)(q-1)}\Big{(}1+\frac{1}{q}(-1)^{n}\Big{)}\cdot(\dim V)^{n}&\text{if $\epsilon$ acts as scalar on $\operatorname{Res}_{G}(V)$}\\ \frac{(p+1)^{r}}{2^{r}(q+1)(q-1)}\cdot(\dim V)^{n}&\text{else.}\end{cases}$
@@
Let $G=\textup{GL}(2,q)$ be the group of units of $kM(2,q)$. By [Kou93, Theorem 4 and Theorem 12], most of the projective indecomposable $kG$-modules are injective when considered as $k\operatorname{M}(2,q)$-modules. So condition ii) in Section 3A is satisfied, and it suffices to compute the asymptotic formula for a faithful $kG$-module. Thus, the statement of the proposition follows from [He25, Theorem 13]. ∎
@@
4C. An example where Section 3A does not apply
Condition ii) in the statement is Section 3A of somewhat restrictive: it is already not satisfied when the monoid is $N=\{1,x,0\}$ where $x^{2}=0$, and of the 35 nonisomorphic monoids of order 4, 9 do not satisfy the condition in characteristic 0. Generalizing the example of $N$, if $M$ is any monoid with group of units $G$, the direct product $M\times N$ (where $N$ may be replaced with any nontrivial nilpotent monoid) will not satisfy condition ii), as any $kG$-module gets a non-split self-extension from the nilpotent element $x$ and so cannot be injective. (These counterexamples were communicated to us by Walter Mazorchuk.)
However, experimentation with monoids of small order (see Section 1B) gives evidence that $V$, when condition i) is satisfied, will still reach $\Gamma^{P}$ even when ii) is not satisfied. As an example, let $M=T_{3}\times N$ where $T_{3}$ is the full transformation monoid on 3 elements and $N=\{1,x,0\}$ as above. There is a unique four-dimensional $kM$-module $V$ satisfying condition i), in characteristic 31. In this case we get $a(n)=2/3\cdot 4^{n}$ as predicted by Equation 1B.1 despite condition ii) not being fulfilled. The fusion graph and action matrix for $V$ are given below, showing that the projective cell is still reached.

### v2

@@ 4. Examples
4.1. Full transformation monoids $T_{n}$, and symmetric inverse monoids $I_{n}$
Let $T_{n}$ and $I_{m}$ denote the full transformation monoid and the symmetric inverse monoid on $m$ elements, respectively. In both cases the group of units is $G=S_{m}$.
Proposition 2.
@@ 4.1. Full transformation monoids $T_{n}$, and symmetric inverse monoids $I_{n}$
(5) $b(n)\sim a(n)=\sum_{z=0}^{\left\lfloor m/2\right\rfloor}\frac{1}{(m-2z)!z!2^{z}}\cdot(\dim V)^{n}=k(n)\sim l(n).$
@@
We know that $kI_{m}$ is semisimple under the assumptions of the theorem (see [Ste16b, Corollary 9.4]). By [Ste16a, Corollary A.3], if $W$ is any $kS_{m}$-module which is not the sign representation, then $\operatorname{Ind}_{G}(W)$ is injective as a $kT_{m}$-module. Thus the assumptions of Theorem 1 are satisfied for both $T_{m}$ and $I_{m}$, and the asymptotic formula for $S_{m}$ given in [CEO24, Example 2.3] now implies the result for $a(n)$. The statement about $k(n)$ follows from Theorem 2. ∎
Remark 6.
One might expect that for $T_{m}$ we can also weaken the requirement that $p=0$ to $p\nmid m!$. However, this is nontrivial to show, as monoid algebras need not have the same structure as we vary between characteristics that do not divide the order of any maximal subgroup. However, for $p$ sufficiently large the result of Proposition 2 will hold. $\Diamond$
Since $kI_{m}$ is semisimple when $p\nmid m!$, the exact formula for $b(n)=l(n)$ can be calculated from Theorem 2. For example, if $V$ is the irreducible $\mathbb{C}I_{3}$-module with character values $0,1,2,0,3,1,0$, then using the character table and decomposition matrix for $\mathbb{C}I_{3}$ which can be found in [Ste16b, Example 9.17], in this case we obtain
@@
We plot below in Figure 2 the fusion graph and action matrix for the 12-dimensional projective $kT_{4}$-module in characteristic 0, and in Figure 3 give the ratio $b(n)/a(n)$ in this case. In this case $\lambda^{\textrm{sec}}=2$, so $\lambda^{\textrm{sec}}/\dim V=1/6$, which explains the rapid convergence. We provide python code in [HT25] to generate the multiplication table of $T_{m}$ and then set up the regular representation of $kT_{m}$ in Magma. As usual, we approximated the characteristic zero behavior by doing the computation in a large enough finite field.
@@
4.2. Matrix monoids $\operatorname{M}(2,q)$
@@
Proposition 3 (Matrix monoids $\operatorname{M}(2,q)$ in nonmodular characteristic).
@@
(6) $a(n)=\begin{cases}\frac{q^{2}+(-1)^{n}}{(q+1)(q^{2}-q)}\cdot(\dim V)^{n}&\text{if $\epsilon$ acts as scalar on $\operatorname{Res}_{G}(V)$},\vskip 5.0pt\\ \frac{q^{2}}{(q+1)(q^{2}-q)}\cdot(\dim V)^{n}&\text{else}.\end{cases}$
@@
Since $\operatorname{M}(2,q)$ is semisimple, by Theorem 1 (or Theorem 2) in this case $a(n)$ is determined by the character table of $G$, which can be found in e.g. [FH91, §5.2]. The sum over the dimensions of irreducible $kG$-modules is $q^{2}(q-1)$, and when $q$ is odd the sum over the column of the character table corresponding to $\epsilon$ is q1q-1. The sum over all other columns in the center vanishes, yielding our formula. ∎
@@
Proposition 4 (Matrix monoids $\operatorname{M}(2,q)$ in defining characteristic).
@@
(7) $a(n)=\begin{cases}\frac{(p+1)^{r}}{2^{r}(q+1)(q-1)}\Big(1+\frac{1}{q}(-1)^{n}\Big)\cdot(\dim V)^{n}&\text{if $\epsilon$ acts as scalar on $\operatorname{Res}_{G}(V)$}\\ \frac{(p+1)^{r}}{2^{r}(q+1)(q-1)}\cdot(\dim V)^{n}&\text{else.}\end{cases}$
@@
Let $G=\textup{GL}(2,q)$ be the group of units of $kM(2,q)$. By [Kou93, Theorem 4 and Theorem 12], most of the projective indecomposable $kG$-modules are injective when considered as $k\operatorname{M}(2,q)$-modules. So condition ii) in Theorem 1 is satisfied, and it suffices to compute the asymptotic formula for a faithful $kG$-module. Thus, the statement of the proposition follows from [He25, Theorem 13]. ∎
@@
4.3. An example where Theorem 1 does not apply
Condition ii) in the statement of Theorem 1 is somewhat restrictive: it is already not satisfied when the monoid is $N=\{1,x,0\}$ where $x^{2}=0$, and of the 35 nonisomorphic monoids of order 4, 9 do not satisfy the condition in characteristic 0. Generalizing the example of $N$, if $M$ is any monoid with group of units $G$, the direct product $M\times N$ (where $N$ may be replaced with any nontrivial nilpotent monoid) will not satisfy condition ii), as any $kG$-module gets a non-split self-extension from the nilpotent element $x$ and so cannot be injective. (These counterexamples were communicated to us by Walter Mazorchuk.)
However, experimentation with monoids of small order (see Section 1.2) gives evidence that $V$, when condition i) is satisfied, will still reach $\Gamma^{P}$ even when ii) is not satisfied. As an example, let $M=T_{3}\times N$ where $T_{3}$ is the full transformation monoid on 3 elements and $N=\{1,x,0\}$ as above. There is a unique four-dimensional $kM$-module $V$ satisfying condition i), in characteristic 31. In this case we get $a(n)=2/3\cdot 4^{n}$ as predicted by Equation 1 despite condition ii) not being fulfilled. The fusion graph and action matrix for $V$ are given below, showing that the projective cell is still reached.